Determine


∫ (x^3 − 2x^2  + 5x) dx


a.     x ^4 / 4 − 2^x3 / 3 + 5x^2 / 2 + c 


b.   x^4 / 4 − 3x^3 / 2 + 2x^2 / 5  + c


c.     x^4 / 4 − 2x^3 / 3 +2x^2 / 5  + c 


d.  x^4 / 4 − 3x^3 / 2 + 5x^2 / 2 + c  



1
Expert's answer
2022-05-02T16:05:37-0400

"\int(x^3 − 2x^2 + 5x) dx =\\ \frac{x^{3+1}}{3+1}-\frac{2x^{2+1}}{2+1}+\frac{5x^{1+1}}{1+1}= \\ \frac{x^{4}}{4}-\frac{2x^{3}}{3}+\frac{5x^{2}}{2}+C"